# Multiple quantum exceptional, diabolical, and hybrid points in multimode bosonic systems: I. Inherited and genuine singularities

### v1

@@ II Two-mode bosonic system: Basic building blocks
Table 1: Real and imaginary parts of the complex eigenfrequencies $\bm{\Lambda_{j}^{\rm r}}-i\bm{\Lambda_{j}^{\rm i}}$ of the matrix $\bm{M^{(2)}}$ given in Eq. (5) for the two-mode bosonic system derived from the equations for the FOMs up to second order. The corresponding moments in the ‘diagonalized’ field operators are written together with their degeneracies (deg.) coming from different positions of the field operators. The DDs of QHPs (partial DDs) derived from the indicated FOMs and the EDs of the constituting QEPs are given. Both genuine and induced QEPs and QHPs are considered. The operators $\bm{\hat{B}_{j}}$ for $j=1,2$ are defined in the lines written for $\bm{\Lambda^{\rm i}_{j}}=\gamma_{+}$ devoted to the first-order FOMs.
III Three-mode bosonic system
The above-analyzed two-mode system provided QHPs with second-order ED and DD. Whereas the DD originates in the mutual coupling between the two modes (described by the strengths $\epsilon$ and $\kappa$), the ED arises from different strengths of damping and amplification of the modes and a given system configuration. Bosonic systems composed of more than two modes give a promise for revealing QHPs with a larger number of EDs and DDs. However, the question is how to choose a suitable configuration of interactions among the modes and their damping and/or amplification rates. Whereas the parity symmetry is useful in seeking promising geometries, the temporal symmetry allows to define suitable relations among the damping and amplification rates. In this section, we begin our investigation by considering three-mode bosonic systems in the linear configuration, and define specific conditions under which QHPs occur. In the following sections, we extend our analysis to the four- and five-mode systems. The geometries together with their characteristic parameters are schematically shown in Fig. 1. We also identify the positions of QHPs in the corresponding parameter spaces. Moreover, we analyze the occurrence of genuine QEPs and QHPs and determine their EDs and DDs from the point of view of the dynamics of second-order FOMs.
Among the bosonic systems with three modes we succeeded in identifying QHPs in the linear configuration that is shown in Fig. 1(b). Similar system in this configuration was experimentally realized in Ref. [59]. We note that a QEP with third-order ED was identified in Ref. [60] in an optomechanical system of three interacting bosonic modes. Hamiltonian $\hat{H}_{3}$ of such a system can be written as follows:
[…]
$\frac{d\hat{\bm{a}}}{dt}$ $=$ $-i\bm{M^{(3)}}\hat{\bm{a}}+\hat{\bm{L}},$ (29)
where
Assuming the condition
$2\gamma_{2}=\gamma_{1}+\gamma_{3},$ (34)
the eigenvalues $\lambda_{j}^{M^{(3)}}$ ($j=1,2,3$) of the dynamical matrix $\bm{M^{(3)}}$ in Eq. (33) share the common damping or amplification rate $\gamma_{+}$ and we may write
$\bm{y_{1}^{M^{(3)}}}$ $=$ $\left[-1,-\frac{i\gamma_{-}}{\xi},1\right]^{T},$
Provided that $\beta=0$, the three eigenvalues $\lambda_{j}^{M^{(3)}}$ for $j=1,2,3$ in Eq. (35), as well as the corresponding eigenvectors in Eq. (36), are equal and, thus, third-order QEPs of the $3\times 3$ matrix $\bm{M^{(3)}}$ are found. Taking into account the structure of the submatrices of the matrix $\bm{M^{(3)}}$, these QEPs are in fact QHPs with second-order DD. Substituting for the eigenvalues $\xi$ of matrix $\bm{\xi}$ from Eq. (13), the condition $\beta=\sqrt{2\zeta^{2}-\gamma_{-}^{2}}=0$ for having a QHP attains the form:
$\frac{\kappa^{2}}{\epsilon^{2}}+\frac{\gamma_{-}^{2}}{2\epsilon^{2}}=1.$ (37)
The analysis of QEPs and QHPs appropriate for the FOM dynamics requires the construction of eigenvalues and eigenvectors of the matrix $\bm{M^{(3)}}$ in its full $6\times 6$ space combining the eigenvalues and eigenvectors in Eqs. (36) and (37) with those of the matrix $\bm{\xi}$ given in Eqs. (13) and (14). They are written in Appendix A. The obtained eigenvectors then allow us to introduce new field operators $\hat{\bm{b}}=\left[\hat{b}_{1},\hat{b}_{1}^{\dagger},\hat{b}_{2},\hat{b}_{3},\hat{b}_{2}^{\dagger},\hat{b}_{3}^{\dagger}\right]^{T}$ in which the dynamical matrix $\bm{M^{(3)}}$ of the Heisenberg-Langevin equations attains its diagonal form $\bm{\Lambda^{(3)}}$, similarly as it was done in Sec. II in the case of the two-mode system. The QEPs and QHPs predicted in the dynamics of the first- and second-order FOMs are summarized in Tab. 2.
@@ III Three-mode bosonic system
$\gamma_{+}$ 0, $\mp\beta$ $\langle b_{1}\rangle$, $\langle b_{2}\rangle$, $\langle b_{2}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{1}}\rangle$ 1 1x3 2x3 1x3 2x3
0, $\mp\beta$ $\langle b_{1}^{\dagger}\rangle$, $\langle b_{3}\rangle$, $\langle b_{3}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{2}}\rangle$ 1 1x3 1x3
$2\gamma_{+}$ 00 $\langle b_{1}^{2}\rangle$ $\langle\bm{\hat{B}_{1}^{2}}\rangle$ 1 1x9 4x9 1x6 2x6
$\mp\beta$ $\langle b_{1}b_{2}\rangle$, $\langle b_{1}b_{2}^{\dagger}\rangle$ 2 +
$\mp 2\beta$ $\langle b_{2}^{2}\rangle$, $\langle b_{2}^{\dagger 2}\rangle$ 1
$\beta-\beta$ $\langle b_{2}^{\dagger}b_{2}\rangle$ 2
0 $\langle b_{1}^{\dagger}b_{1}\rangle$ $\langle\bm{\hat{B}_{2}\hat{B}_{1}}\rangle$ 2 2x9 1x9 1x9
$\mp\beta$ $\langle b_{1}^{\dagger}b_{2}\rangle$, $\langle b_{1}^{\dagger}b_{2}^{\dagger}\rangle$ 2
$\mp\beta$ $\langle b_{3}b_{1}\rangle$, $\langle b_{3}^{\dagger}b_{1}\rangle$ 2
$\mp 2\beta$ $\langle b_{3}b_{2}\rangle$, $\langle b_{3}^{\dagger}b_{2}^{\dagger}\rangle$ 2
$\beta-\beta$ $\langle b_{3}b_{2}^{\dagger}\rangle$, $\langle b_{3}^{\dagger}b_{2}\rangle$ 2
00 $\langle b_{1}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{2}^{2}}\rangle$ 1 1x9 1x6
$\mp\beta$ $\langle b_{1}^{\dagger}b_{3}\rangle$, $\langle b_{1}^{\dagger}b_{3}^{\dagger}\rangle$ 2
$\mp 2\beta$ $\langle b_{3}^{2}\rangle$, $\langle b_{3}^{\dagger 3}\rangle$ 1
$\beta-\beta$ $\langle b_{3}^{\dagger}b_{3}\rangle$ 2
@@ 00 $\langle b_{1}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{2}^{2}}\rangle$ 1 1x9 1x6
Table 2: Real and imaginary parts of the complex eigenfrequencies $\bm{\Lambda_{j}^{\rm r}}-i\bm{\Lambda_{j}^{\rm i}}$ of the matrix $\bm{M^{(3)}}$, given in Eq. (33), for the linear three-mode bosonic system derived from the equations for the FOMs up to second order. For details, see the caption to Tab. 1.
IV Four-mode bosonic systems
IV.1 Linear configurations
$\hat{H}_{4,{\rm l}}$ $=$ $\Bigl{[}\hbar\epsilon\hat{a}_{1}^{\dagger}\hat{a}_{2}+\hbar\epsilon\hat{a}_{2}^{\dagger}\hat{a}_{3}+\hbar\epsilon\hat{a}_{3}^{\dagger}\hat{a}_{4}+\hbar\kappa\hat{a}_{1}\hat{a}_{2}$ (38)
[…]
$\frac{d\hat{\bm{a}}}{dt}$ $=$ $-i\bm{M^{(4)}_{\rm l}}\hat{\bm{a}}+\hat{\bm{L}},$ (39)
Using the $2\times 2$ submatrices defined in Eq. (10) we write the dynamical matrix $\bm{M^{(4)}_{\rm l}}$:
QEPs and QHPs have been found in the following two configurations:
IV.1.1 Linear configuration with equal damping and/or amplification rates of neighbor modes
In this case, diagonalization of the dynamical matrix $\bm{M^{(4)}_{\rm l1}}$ in Eq. (44) provides the eigenvalues
$\lambda_{1,2}^{M^{(4)}_{\rm l1}}$ $=$ $-i\gamma_{+}\pm{\alpha}_{-},$
$\lambda_{3,4}^{M^{(4)}_{\rm l1}}$ $=$ $-i\gamma_{+}\pm{\alpha}_{+},$ (46)
and the corresponding eigenvectors
If $\mu=0$ then ${\alpha}_{+}={\alpha}_{-}$ and $\chi_{+}=\chi_{-}$. The eigenvalues in Eq. (46) are doubly degenerated: $\lambda_{1}^{M^{(4)}_{\rm l1}}=\lambda_{3}^{M^{(4)}_{\rm l1}}$ and $\lambda_{2}^{M^{(4)}_{\rm l1}}=\lambda_{4}^{M^{(4)}_{\rm l1}}$. The same holds also for their eigenvectors given in Eq. (47) and we have two QEPs with second-order ED. Similarly as in the case of three-mode bosonic system, the inclusion of the structure of submatrices in the matrix $\bm{M^{(4)}_{\rm l1}}$ gives the second-order DD to these QEPs and we, thus, have two QHPs. They are localized at the ellipse
$\frac{\kappa^{2}}{\epsilon^{2}}+\frac{16\gamma_{-}^{2}}{5\epsilon^{2}}=1$ (48)
defined in the parameter space $(\kappa/\epsilon,\gamma_{-}/\epsilon)$. Real parts of the eigenvalues $\lambda_{j}^{M^{(4)}_{\rm l1}}$ for $j=1,\ldots,4$ are drawn in this space in Fig. 2(c).
The eigenvalues and eigenvectors of the $8\times 8$ matrix $\bm{M^{(4)}_{\rm l1}}$ obtained by merging those written in Eqs. (46) and (47) and Eqs. (13) and (14) for the matrix $\bm{\xi}$ (for details, see Appendix A) define the transformation into the new field operators $\hat{\bm{b}}=\left[\hat{b}_{1},\hat{b}_{2},\hat{b}_{1}^{\dagger},\hat{b}_{2}^{\dagger},\hat{b}_{3},\hat{b}_{4},\hat{b}_{3}^{\dagger},\hat{b}_{4}^{\dagger}\right]^{T}$ in which the Heisenberg-Langevin equations have the diagonal form. This allows to reveal QEPs and QHPs in the dynamics of the first- and second-order FOMs. They are summarized in Tab. 3.
@@ IV.1.1 Linear configuration with equal damping and/or amplification rates of neighbor modes
$\gamma_{+}$ $\alpha_{\pm}$ $\langle b_{1}\rangle$, $\langle b_{3}\rangle$ $\langle\bm{\hat{B}_{1}}\rangle$ 1 2x2 4x2 2x2 4x2
$\langle b_{4}\rangle$, $\langle b_{2}\rangle$ $\langle\bm{\hat{B}_{2}}\rangle$ 1
$-\alpha_{\pm}$ $\langle b_{1}^{\dagger}\rangle$, $\langle b_{3}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{3}}\rangle$ 1 2x2 2x2
$\langle b_{4}^{\dagger}\rangle$, $\langle b_{2}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{4}}\rangle$ 1
$2\gamma_{+}$ $2\alpha_{\pm}$ $\langle b_{1}^{2}\rangle$, $\langle b_{3}^{2}\rangle$ $\langle\bm{\hat{B}_{1}^{2}}\rangle$ 1 1x4 16x4 1x3 4x3
$\alpha_{-}+\alpha_{+}$ $\langle b_{1}b_{3}\rangle$ 2 +
$2\alpha_{\mp}$ $\langle b_{2}^{2}\rangle$, $\langle b_{4}^{2}\rangle$ $\langle\bm{\hat{B}_{2}^{2}}\rangle$ 1 1x4 1x3
$\alpha_{-}+\alpha_{+}$ $\langle b_{2}b_{4}\rangle$ 2
$-2\alpha_{\pm}$ $\langle b_{1}^{\dagger 2}\rangle$, $\langle b_{3}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{3}^{2}}\rangle$ 1 1x4 1x3
$-\alpha_{-}-\alpha_{+}$ $\langle b_{1}^{\dagger}b_{3}^{\dagger}\rangle$ 2
$-2\alpha_{\mp}$ $\langle b_{2}^{\dagger 2}\rangle$, $\langle b_{4}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{4}^{2}}\rangle$ 1 1x4 1x3
$-\alpha_{-}-\alpha_{+}$ $\langle b_{2}^{\dagger}b_{4}^{\dagger}\rangle$ 2
$-\alpha_{\pm}+\alpha_{\pm}$ $\langle b_{1}^{\dagger}b_{1}\rangle$, $\langle b_{3}^{\dagger}b_{3}\rangle$ $\langle\bm{\hat{B}_{3}\hat{B}_{1}}\rangle$ 2 2x4 1x4 6x4
±$\pm\alpha_{-}\mp\alpha_{+}$ $\langle b_{1}^{\dagger}b_{3}\rangle$, $\langle b_{1}b_{3}^{\dagger}\rangle$ 2
$-\alpha_{\mp}+\alpha_{\mp}$ $\langle b_{2}^{\dagger}b_{2}\rangle$, $\langle b_{4}^{\dagger}b_{4}\rangle$ $\langle\bm{\hat{B}_{4}\hat{B}_{2}}\rangle$ 2 2x4 1x4
$\mp\alpha_{-}\pm\alpha_{+}$ $\langle b_{2}^{\dagger}b_{4}\rangle$, $\langle b_{4}^{\dagger}b_{2}\rangle$ 2
$\alpha_{+}+\alpha_{-}$ $\langle b_{1}b_{2}\rangle$, $\langle b_{3}b_{4}\rangle$ $\langle\bm{\hat{B}_{1}\hat{B}_{2}}\rangle$ 2 2x4 1x4
$2\alpha_{\pm}$ $\langle b_{1}b_{4}\rangle$, $\langle b_{3}b_{2}\rangle$ 2
$\alpha_{\mp}-\alpha_{\pm}$ $\langle b_{1}^{\dagger}b_{2}\rangle$, $\langle b_{3}^{\dagger}b_{4}\rangle$ $\langle\bm{\hat{B}_{3}\hat{B}_{2}}\rangle$ 2 2x4 1x4
$\alpha_{\pm}-\alpha_{\pm}$ $\langle b_{1}^{\dagger}b_{4}\rangle$, $\langle b_{3}^{\dagger}b_{2}\rangle$ 2
$\alpha_{\pm}-\alpha_{\mp}$ $\langle b_{2}^{\dagger}b_{1}\rangle$, $\langle b_{4}^{\dagger}b_{3}\rangle$ $\langle\bm{\hat{B}_{4}\hat{B}_{1}}\rangle$ 2 2x4 1x4
$\alpha_{\pm}-\alpha_{\pm}$ $\langle b_{4}^{\dagger}b_{1}\rangle$, $\langle b_{2}^{\dagger}b_{3}\rangle$ 2
$-\alpha_{+}-\alpha_{-}$ $\langle b_{1}^{\dagger}b_{2}^{\dagger}\rangle$, $\langle b_{3}^{\dagger}b_{4}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{3}\hat{B}_{4}}\rangle$ 2 2x4 1x4
$-2\alpha_{\pm}$ $\langle b_{1}^{\dagger}b_{4}^{\dagger}\rangle$, $\langle b_{3}^{\dagger}b_{2}^{\dagger}\rangle$ 2
@@
Table 3: Real and imaginary parts of the complex eigenfrequencies $\bm{\Lambda_{j}^{\rm r}}-i\bm{\Lambda_{j}^{\rm i}}$ of the matrix $\bm{M^{(4)}_{\rm l1}}$, given in Eq. (44) with Eq. (45), for the linear four-mode bosonic system with equal damping and/or amplification rates for neighbor modes derived from the equations for the FOMs up to second order. We note that ${\alpha}_{\pm}\left(\zeta\right)={\alpha}_{\mp}\left(-\zeta\right)$ is used here. For details, see the caption to Tab. 1.
IV.1.2 Linear configuration with different damping and/or amplification rates for neighbor modes
We assume equal the damping and/or amplification rates in modes 1 and 3 and also in modes 2 and 4:
$\lambda_{1,2}^{M^{(4)}_{\rm l2}}$ $=$ $-i\gamma_{+}\pm\alpha_{+},$
$\lambda_{3,4}^{M^{(4)}_{\rm l2}}$ $=$ $-i\gamma_{+}\pm\alpha_{-}.$ (50)
The corresponding eigenvectors are reached in the form:
Provided that $\alpha_{+}=0$ [$\alpha_{-}=0$] the eigenvalues $\lambda_{1}^{M^{(4)}_{\rm l2}}$ and $\lambda_{2}^{M^{(4)}_{\rm l2}}$ [$\lambda_{3}^{M^{(4)}_{\rm l2}}$ and $\lambda_{4}^{M^{(4)}_{\rm l2}}$] in Eq. (50) and the eigenvectors $\bm{y_{1}^{M^{(4)}_{\rm l2}}}$ and $\bm{y_{2}^{M^{(4)}_{\rm l2}}}$ [$\bm{y_{3}^{M^{(4)}_{\rm l2}}}$ and $\bm{y_{4}^{M^{(4)}_{\rm l2}}}$] in Eq. (51) are equal to each other and we have a QEP with second-order ED. This means that, for the $8\times 8$ dynamical matrix $\bm{M^{(4)}_{\rm l2}}$, we predict a QHP with second-order ED and DD observed either for $\alpha_{+}=0$ or $\alpha_{-}=0$. These conditions define two ellipses in the parameter space $(\kappa/\epsilon,\gamma_{-}/\epsilon)$,
They are shown in Fig. 2(d) where the real parts of the eigenvalues $\lambda_{j}^{M^{(4)}_{\rm l2}}$ for $j=1,\ldots,4$ are plotted.
The eigenvalues and eigenvectors of the $8\times 8$ matrix $\bm{M^{(4)}_{\rm l2}}$, given in Appendix A, determine the transformation that leaves the Heisenberg-Langevin equations in their diagonal form. Thus, we define the new field operators $\hat{\bm{b}}=\left[\hat{b}_{1},\hat{b}_{2},\hat{b}_{1}^{\dagger},\hat{b}_{2}^{\dagger},\hat{b}_{3},\hat{b}_{4},\hat{b}_{3}^{\dagger},\hat{b}_{4}^{\dagger}\right]^{T}$, whose evolution identifies the QEPs and QHPs in the first- and second-order FOM dynamics given in Tab. 4.

### v3

@@ Appendix B QEPs and QHPs in first- and second-order field-operator-moments spaces: Comparison
Table 1: Real and imaginary parts of the complex eigenfrequencies $\bm{\Lambda_{j}^{\rm r}}-i\bm{\Lambda_{j}^{\rm i}}$ of the matrix $\bm{M^{(2)}}$ given in Eq. (5) for the two-mode bosonic system derived from the equations for the FOMs up to second order. The corresponding moments in the ‘diagonalized’ field operators are written together with their degeneracies (deg.) coming from different positions of the field operators. The DDs of QHPs (partial $D$s) derived from the indicated FOMs and the EDs of the constituting QEPs are given. Both genuine and induced QEPs and QHPs are considered. The operator vectors $\bm{\hat{B}_{j}}$ for $j=1,2$ are defined in the rows written for $\bm{\Lambda^{\rm i}_{j}}=\gamma_{+}$ devoted to the first-order FOMs, i.e. $\bm{\hat{B}_{j}}\equiv[\hat{b}_{j},\hat{b}_{j}^{\dagger}]$ for $j=1,2$. Symbol $\bm{\hat{B}_{j}\hat{B}_{k}}$, $j,k=1,2$, stands for the tensorial product that provides four terms explicitly written in the rows for $\bm{\Lambda^{\rm i}_{j}}=2\gamma_{+}$; the terms derived from those explicitly written by using the commutation relations are omitted.
Considering the linear three-mode bosonic system with its dynamical matrix ${\bm{M}}^{(3)}$ given in Eq. (38) under the condition given in Eq. (39), the QEPs and QHPs predicted in the dynamics of first- and second-order FOMs are summarized in Tab. 2. As shown there, the inherited QEPs with a third-order ED give rise to genuine QEPs with a ninth-order ED when analyzing the dynamics of second-order FOMs.
@@
$\gamma_{+}$ 0, $\mp\beta$ $\langle\hat{b}_{1}\rangle$, $\langle\hat{b}_{2}\rangle$, $\langle\hat{b}_{2}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{1}}\rangle$ 1 1x3 2x3 1x3 2x3
0, $\mp\beta$ $\langle\hat{b}_{1}^{\dagger}\rangle$, $\langle\hat{b}_{3}\rangle$, $\langle\hat{b}_{3}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{2}}\rangle$ 1 1x3 1x3
$2\gamma_{+}$ 0 $\langle\hat{b}_{1}^{2}\rangle$ $\langle\bm{\hat{B}_{1}^{2}}\rangle$ 1 1x9 4x9 1x6 2x6
$\mp\beta$ $\langle\hat{b}_{1}\hat{b}_{2}\rangle$, $\langle\hat{b}_{1}\hat{b}_{2}^{\dagger}\rangle$ 2 +
$\mp 2\beta$ $\langle\hat{b}_{2}^{2}\rangle$, $\langle\hat{b}_{2}^{\dagger 2}\rangle$ 1
$\beta-\beta$ $\langle\hat{b}_{2}^{\dagger}\hat{b}_{2}\rangle$ 2
0 $\langle\hat{b}_{1}^{\dagger}\hat{b}_{1}\rangle$ $\langle\bm{\hat{B}_{2}\hat{B}_{1}}\rangle$ 2 2x9 1x9 1x9
$\mp\beta$ $\langle\hat{b}_{1}^{\dagger}\hat{b}_{2}\rangle$, $\langle\hat{b}_{1}^{\dagger}\hat{b}_{2}^{\dagger}\rangle$ 2
$\mp\beta$ $\langle\hat{b}_{3}\hat{b}_{1}\rangle$, $\langle\hat{b}_{3}^{\dagger}\hat{b}_{1}\rangle$ 2
$\mp 2\beta$ $\langle\hat{b}_{3}\hat{b}_{2}\rangle$, $\langle\hat{b}_{3}^{\dagger}\hat{b}_{2}^{\dagger}\rangle$ 2
$\beta-\beta$ $\langle\hat{b}_{3}\hat{b}_{2}^{\dagger}\rangle$, $\langle\hat{b}_{3}^{\dagger}\hat{b}_{2}\rangle$ 2
0 $\langle\hat{b}_{1}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{2}^{2}}\rangle$ 1 1x9 1x6
$\mp\beta$ $\langle\hat{b}_{1}^{\dagger}\hat{b}_{3}\rangle$, $\langle\hat{b}_{1}^{\dagger}\hat{b}_{3}^{\dagger}\rangle$ 2
$\mp 2\beta$ $\langle\hat{b}_{3}^{2}\rangle$, $\langle\hat{b}_{3}^{\dagger 3}\rangle$ 1
$\beta-\beta$ $\langle\hat{b}_{3}^{\dagger}\hat{b}_{3}\rangle$ 2
@@ 0 $\langle\hat{b}_{1}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{2}^{2}}\rangle$ 1 1x9 1x6
Table 2: Real and imaginary parts of the complex eigenfrequencies $\bm{\Lambda_{j}^{\rm r}}-i\bm{\Lambda_{j}^{\rm i}}$ of the matrix $\bm{M^{(3)}}$, given in Eq. (38), for the linear three-mode bosonic system derived from the equations for the FOMs up to second order. We have $\bm{\hat{B}_{1}}\equiv[\hat{b}_{1},\hat{b}_{2},\hat{b}_{2}^{\dagger}]$ and $\bm{\hat{B}_{2}}\equiv[\hat{b}_{1}^{\dagger},\hat{b}_{3},\hat{b}_{3}^{\dagger}]$ and more details are given in the caption to Tab. 1.
In the linear four-mode bosonic system described by the dynamical matrix ${\bm{M}}_{\rm l}^{(4)}$ given in Eq. (49), which features equal damping and/or amplification rates for neighboring modes [see Eq. (50)], the QEPs and QHPs in the dynamics of first- and second-order FOMs are summarized in Tab. 3. These results also apply to the circular four-mode bosonic system with dynamical matrix ${\bm{M}}_{\rm c}^{(4)}$ shown in Eq. (63), assuming equal damping and/or amplification rates of neighboring modes [see Eq. (64]. According to Tab. 3, the inherited QEPs with doubled second-order ED give rise to genuine QEPs exhibiting fourth-order ED and sixth-order DD when analyzing second-order FOM dynamics.
@@
$\gamma_{+}$ $\alpha_{\pm}$ $\langle\hat{b}_{1}\rangle$, $\langle\hat{b}_{3}\rangle$ $\langle\bm{\hat{B}_{1}}\rangle$ 1 2x2 4x2 2x2 4x2
$\langle\hat{b}_{4}\rangle$, $\langle\hat{b}_{2}\rangle$ $\langle\bm{\hat{B}_{2}}\rangle$ 1
$\alpha_{\pm}$ $\langle\hat{b}_{1}^{\dagger}\rangle$, $\langle\hat{b}_{3}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{3}}\rangle$ 1 2x2 2x2
$\langle\hat{b}_{4}^{\dagger}\rangle$, $\langle\hat{b}_{2}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{4}}\rangle$ 1
$2\gamma_{+}$ $2\alpha_{\pm}$ $\langle\hat{b}_{1}^{2}\rangle$, $\langle\hat{b}_{3}^{2}\rangle$ $\langle\bm{\hat{B}_{1}^{2}}\rangle$ 1 1x4 16x4 1x3 4x3
$\alpha_{-}+\alpha_{+}$ $\langle\hat{b}_{1}\hat{b}_{3}\rangle$ 2 +
$2\alpha_{\mp}$ $\langle\hat{b}_{2}^{2}\rangle$, $\langle\hat{b}_{4}^{2}\rangle$ $\langle\bm{\hat{B}_{2}^{2}}\rangle$ 1 1x4 1x3
$\alpha_{-}+\alpha_{+}$ $\langle\hat{b}_{2}\hat{b}_{4}\rangle$ 2
$2\alpha_{\pm}$ $\langle\hat{b}_{1}^{\dagger 2}\rangle$, $\langle\hat{b}_{3}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{3}^{2}}\rangle$ 1 1x4 1x3
$\alpha_{-}-\alpha_{+}$ $\langle\hat{b}_{1}^{\dagger}\hat{b}_{3}^{\dagger}\rangle$ 2
$2\alpha_{\mp}$ $\langle\hat{b}_{2}^{\dagger 2}\rangle$, $\langle\hat{b}_{4}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{4}^{2}}\rangle$ 1 1x4 1x3
$\alpha_{-}-\alpha_{+}$ $\langle\hat{b}_{2}^{\dagger}\hat{b}_{4}^{\dagger}\rangle$ 2
$\alpha_{\pm}+\alpha_{\pm}$ $\langle\hat{b}_{1}^{\dagger}\hat{b}_{1}\rangle$, $\langle\hat{b}_{3}^{\dagger}\hat{b}_{3}\rangle$ $\langle\bm{\hat{B}_{3}\hat{B}_{1}}\rangle$ 2 2x4 1x4 6x4
$\pm\alpha_{-}\mp\alpha_{+}$ $\langle\hat{b}_{1}^{\dagger}\hat{b}_{3}\rangle$, $\langle\hat{b}_{1}\hat{b}_{3}^{\dagger}\rangle$ 2
$\alpha_{\mp}+\alpha_{\mp}$ $\langle\hat{b}_{2}^{\dagger}\hat{b}_{2}\rangle$, $\langle\hat{b}_{4}^{\dagger}\hat{b}_{4}\rangle$ $\langle\bm{\hat{B}_{4}\hat{B}_{2}}\rangle$ 2 2x4 1x4
$\mp\alpha_{-}\pm\alpha_{+}$ $\langle\hat{b}_{2}^{\dagger}\hat{b}_{4}\rangle$, $\langle\hat{b}_{4}^{\dagger}\hat{b}_{2}\rangle$ 2
$\alpha_{+}+\alpha_{-}$ $\langle\hat{b}_{1}\hat{b}_{2}\rangle$, $\langle\hat{b}_{3}\hat{b}_{4}\rangle$ $\langle\bm{\hat{B}_{1}\hat{B}_{2}}\rangle$ 2 2x4 1x4
$2\alpha_{\pm}$ $\langle\hat{b}_{1}\hat{b}_{4}\rangle$, $\langle\hat{b}_{3}\hat{b}_{2}\rangle$ 2
$\alpha_{\mp}-\alpha_{\pm}$ $\langle\hat{b}_{1}^{\dagger}\hat{b}_{2}\rangle$, $\langle\hat{b}_{3}^{\dagger}\hat{b}_{4}\rangle$ $\langle\bm{\hat{B}_{3}\hat{B}_{2}}\rangle$ 2 2x4 1x4
$\alpha_{\pm}-\alpha_{\pm}$ $\langle\hat{b}_{1}^{\dagger}\hat{b}_{4}\rangle$, $\langle\hat{b}_{3}^{\dagger}\hat{b}_{2}\rangle$ 2
$\alpha_{\pm}-\alpha_{\mp}$ $\langle\hat{b}_{2}^{\dagger}\hat{b}_{1}\rangle$, $\langle\hat{b}_{4}^{\dagger}\hat{b}_{3}\rangle$ $\langle\bm{\hat{B}_{4}\hat{B}_{1}}\rangle$ 2 2x4 1x4
$\alpha_{\pm}-\alpha_{\pm}$ $\langle\hat{b}_{4}^{\dagger}\hat{b}_{1}\rangle$, $\langle\hat{b}_{2}^{\dagger}\hat{b}_{3}\rangle$ 2
$\alpha_{+}-\alpha_{-}$ $\langle\hat{b}_{1}^{\dagger}\hat{b}_{2}^{\dagger}\rangle$, $\langle\hat{b}_{3}^{\dagger}\hat{b}_{4}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{3}\hat{B}_{4}}\rangle$ 2 2x4 1x4
$2\alpha_{\pm}$ $\langle\hat{b}_{1}^{\dagger}\hat{b}_{4}^{\dagger}\rangle$, $\langle\hat{b}_{3}^{\dagger}\hat{b}_{2}^{\dagger}\rangle$ 2
@@
Table 3: Real and imaginary parts of the complex eigenfrequencies $\bm{\Lambda_{j}^{\rm r}}-i\bm{\Lambda_{j}^{\rm i}}$ of the matrix $\bm{M^{(4)}_{\rm l1}}$, given in Eq. (49) with Eq. (50), for the linear four-mode bosonic system with equal damping and/or amplification rates for neighbor modes derived from the equations for the FOMs up to second order. We note that ${\alpha}_{\pm}\left(\zeta\right)={\alpha}_{\mp}\left(-\zeta\right)$ is used here. We have $\bm{\hat{B}_{1}}\equiv[\hat{b}_{1},\hat{b}_{3}]$, $\bm{\hat{B}_{2}}\equiv[\hat{b}_{2},\hat{b}_{4}]$, $\bm{\hat{B}_{3}}=\bm{\hat{B}_{1}}^{\dagger}$, $\bm{\hat{B}_{4}}=\bm{\hat{B}_{2}}^{\dagger}$. More details are given in the caption to Tab. 1.
On the other hand, for the linear four-mode bosonic system described by the dynamical matrix ${\bm{M}}_{\rm l}^{(4)}$ given in Eq. (49), with varying damping and/or amplification rates of neighboring modes [see Eq. (54)], the QEPs and QHPs observed in the dynamics of first- and second-order FOMs are summarized in Tab. 4. The spectral structure of first-order FOMs, featuring two inherited QEPs with second-order EDs, evolves into a more complex second-order FOM structure that includes genuine QEPs with second-, third-, and fourth-order EDs. Notably, the genuine QHPs with second-order ED exhibit a sixteenth-order DD.